# New gravitational wave polarization modes in the torsionless spacetime

### v2

@@ Appendix A Derivation of the geodesic equation in general relativity
^$\widehat{\nabla}_{\mu}T^{\mu\nu}=\frac{1}{\sqrt{-g}}\partial_{\mu}\left(\sqrt{-g}T^{\mu\nu}\right)+\widehat{\Gamma}^{\nu}_{~{}\mu\lambda}T^{\mu\lambda}=0.$ (70)
@@ Appendix B Proof of the equivalence of assumptions
[…]
@@ Appendix C Construction of the second-order perturbation action for the most general Palatini theory
$+$ $\!\frac{1}{2}D_{(1)}\eta^{\rho\nu}\delta^{\mu}_{~{}\lambda}\left(\partial_{\lambda}h\right)+\frac{1}{2}D_{(1)}\eta^{\rho\mu}\delta^{\nu}_{~{}\lambda}\left(\partial_{\lambda}h\right)+E_{(1)}\eta^{\rho\sigma}\eta^{\mu\nu}\left(\partial_{\rho}h_{\lambda\sigma}\right)$
@@
$=$ $\!0$
Note that linear perturbation equations (101) and (102) have the form of “tensor$=0$”, so we can uniquely decompose them in the same way as Eq. (II). Owing to the uniqueness of the decomposition, the spatial tensor, vector, and scalar components in the linear perturbation equations must each independently equal 00. By substituting Eq. (II) into these equations and expressing them in terms of gauge invariants via Eqs. (10)-(13), we find that the spatial tensor, vector, and scalar equations depend exclusively on the gauge-invariant spatial tensors, vectors, and scalars. Thus, the equations have been decoupled. This enables independent consideration of tensor, vector, and scalar perturbations, corresponding to the case of tensor, vector, and scalar mode gravitational waves. Given the large number and complexity of these equations, their explicit forms are provided in Appendix E. These decoupled equations are not fully independent; the number of independent equations is the same as the number of independent gauge invariant perturbations (10)-(13). Solving these decoupled equations forms the basis for analyzing the polarization modes of gravitational waves in.
@@ D.1 Case 1: homogeneous field equations
In this case, if the test particles do not possess hypermomentum, then the presence of $\Sigma^{\lambda}_{~{}\mu\nu}$ will have no impact on the problem under consideration, and the situation will revert to general relativity. When the particles possess hypermomentum, although the $\Sigma^{\lambda}_{~{}\mu\nu}$ does not influence the motion of $h_{\mu\nu}$, it still impacts the relative motion between the test particles, thereby affects the polarization modes of gravitational waves. However, assuming the existence of a hypermomentum charge in this case would result in physical inconsistency. By noting that the gauge invariant $K_{i}$ in Eq. (II) is composed of $h_{\mu\nu}$ and $\Sigma^{\lambda}_{~{}\mu\nu}$, we can conclude that the gauge-invariant field equations in this case impose no constraints on $K_{i}$. However, as indicated by Eq. (IV), $K_{i}$ can lead to observable polarization effects, making this situation unreasonable.
@@ D.2 Case 2: the connection as a constraint variable
From this equation and Eq. (IV), it follows that tensor modes—specifically, the $+$ and $\times$ modes—propagate at the speed of light.
@@
$-\frac{1}{2}\left(A_{(1)}-4A_{(2)}-C_{(1)}\right)\left(2\phi+\Theta\right)\!$ $=$ $\!0$ (109)
@@ D.3 Case 3: the metric as a constraint variable
$0\displaystyle 2A_{(2)}-\frac{1}{2}A_{(1)}+\frac{1}{2}C_{(1)}-B_{(1)}=0,~{}-2A_{(2)}-\frac{1}{2}C_{(1)}-\frac{1}{2}E_{(1)}=$ 0 . (112)
@@
Here, $\mathring{h}^{TT}_{ij}$ and $\mathring{S}_{ij}$ are constants. By substituting these plane wave forms into the decoupled equations, we can transform the system of differential equations into a linear algebraic system (system of linear equations), due to the linearity of the equations and their constant coefficients. At this stage, the possible wave vector $k^{\mu}$ can be determined by the condition that the determinant of the coefficient matrix is zero, yielding the dispersion relation of gravitational waves. For the corresponding $k^{\mu}$, solving the linear system provides the relationship between gauge invariant perturbations and determines the polarization modes of gravitational waves. Additionally, due to spatial rotational symmetry, we can assume that gravitational waves propagate in the $+z$ direction without loss of generality, simplifying the equations to be solved. Henceforth, we assume gravitational waves propagate along the $+z$ direction.
@@
According to Eq. (IV), only $h^{TT}_{ij}$ contributes to the polarization modes. However, in the solutions corresponding to $m_{2}$ and $m_{3}$ in Eq. (116), one can obtian $h^{TT}_{ij}=0$, which means these solutions do not excite tensor mode gravitational waves. Therefore, tensor modes, specifically the $+$ and $\times$ modes, can only propagate as massless modes at the speed of light or propagate with mass $m_{4}$. It is important to note that tensor mode gravitational waves, propagating at the speed of light, always exist.
@@ D.4 Case 4: no constraint variables
Here, the expressions for $\mathcal{A}$, $\mathcal{B}$, $\mathcal{C}$, and $\mathcal{D}$ are exceptionally complex, so they will not be listed here. Similarly, in the solutions corresponding to $m_{2}$ and $m_{3}$, $h^{TT}_{ij}=0$, so it does not actually contribute to tensor mode gravitational waves. The $+$ and $\times$ modes that propagate at the speed of light always exist. The difference between the vector and scalar modes and Case 3 is that they possess more degrees of freedom. For the same reasons, a general analysis remains challenging.
@@ Appendix E Decoupled equations
$-\left(2A_{(1)}+2B_{(1)}+\frac{1}{2}C_{(1)}+E_{(1)}\right)N_{i}-\left(2B_{(1)}+4E_{(1)}\right)h_{i}-\left(8B_{(1)}+6E_{(1)}\right)q_{i}$
@@
$\left(A_{(1)}+\frac{1}{2}C_{(1)}\right)C_{i}-\left(A_{(1)}+\frac{1}{2}C_{(1)}\right)\bar{C}_{i}-\frac{1}{2}C_{(1)}K_{i}-\frac{1}{2}C_{(1)}\left(K_{i}-\Xi_{i}\right)$ (133)
@@
$\left.+2B_{(2)}\partial_{0}h_{i}+2B_{(2)}\partial_{0}N_{i}+8B_{(2)}\partial_{0}q_{i}+B_{(2)}\partial_{0}\Delta f_{i}\right)=0,$

### v3

@@ Appendix A Derivation of the geodesic equation in general relativity
$\widehat{\nabla}_{\mu}T^{\mu\nu}=\frac{1}{\sqrt{-g}}\partial_{\mu}\left(\sqrt{-g}T^{\mu\nu}\right)+\widehat{\Gamma}^{\nu}_{~\mu\lambda}T^{\mu\lambda}=0.$ (71)
@@ Appendix B Proof of the equivalence of assumptions
$=$ $\!\!\sum_{l,m,n\in\mathbb{N}_{0}}a_{lmn}\left(\int dxe^{\lambda_{1}x}\delta^{(l)}(x)\right)\left(\int dye^{\lambda_{2}y}\delta^{(m)}(y)\right)\left(\int dze^{\lambda_{3}z}\delta^{(n)}(z)\right)$
@@ Appendix C Construction of the second-order perturbation action for the most general Palatini theory
$+$ $\!\frac{1}{2}D_{(1)}\eta^{\rho\nu}\delta^{\mu}_{~\lambda}\left(\partial_{\rho}h\right)+\frac{1}{2}D_{(1)}\eta^{\rho\mu}\delta^{\nu}_{~\lambda}\left(\partial_{\rho}h\right)+E_{(1)}\eta^{\rho\sigma}\eta^{\mu\nu}\left(\partial_{\rho}h_{\lambda\sigma}\right)$
@@
$=$ $\!0.$
Note that linear perturbation equations (102) and (103) have the form of “tensor=0=0”, so we can uniquely decompose them in the same way as Eq. (II). Owing to the uniqueness of the decomposition, the spatial tensor, vector, and scalar components in the linear perturbation equations must each independently equal 0. By substituting Eq. (II) into these equations and expressing them in terms of gauge invariants via Eqs. (10)-(13), we find that the spatial tensor, vector, and scalar equations depend exclusively on the gauge-invariant spatial tensors, vectors, and scalars. Thus, the equations have been decoupled. This enables independent consideration of tensor, vector, and scalar perturbations, corresponding to the case of tensor, vector, and scalar mode gravitational waves. Given the large number and complexity of these equations, their explicit forms are provided in Appendix E. These decoupled equations are not fully independent; the number of independent equations is the same as the number of independent gauge invariant perturbations (10)-(13). Solving these decoupled equations forms the basis for analyzing the polarization modes of gravitational waves in.
@@ D.1 Case 1: homogeneous field equations
In this case, if the test particles do not possess hypermomentum (or the effect of hypermomentum is negligible), then the presence of $\Sigma^{\lambda}_{~\mu\nu}$ will have no impact on the problem under consideration, and the situation will revert to general relativity. When the particles possess hypermomentum, although the $\Sigma^{\lambda}_{~\mu\nu}$ does not influence the motion of $h_{\mu\nu}$, it still impacts the relative motion between the test particles, thereby affects the polarization modes of gravitational waves. However, assuming the existence of a hypermomentum charge in this case would result in physical inconsistency. By noting that the gauge invariant $K_{i}$ in Eq. (II) is composed of $h_{\mu\nu}$ and $\Sigma^{\lambda}_{~\mu\nu}$, we can conclude that the gauge-invariant field equations in this case impose no constraints on $K_{i}$. However, as indicated by Eq. (IV), $K_{i}$ can lead to observable polarization effects, making this situation unreasonable.
@@ D.2 Case 2: the connection as a constraint variable
From this equation and Eq. (IV), it follows that tensor modes—specifically, the ++ and $\times$ modes—propagate at the speed of light.
@@
$-\frac{1}{2}\left(A_{(1)}-4A_{(2)}-C_{(1)}\right)\left(2\phi+\Theta\right)\!$ $=$ $\!0.$ (110)
@@ D.3 Case 3: the metric as a constraint variable
$2A_{(2)}-\frac{1}{2}A_{(1)}+\frac{1}{2}C_{(1)}-B_{(1)}=0,~-2A_{(2)}-\frac{1}{2}C_{(1)}-\frac{1}{2}E_{(1)}=0.$ (113)
@@
Here, $\mathring{h}^{TT}_{ij}$ and $\mathring{S}_{ij}$ are constants. By substituting these plane wave forms into the decoupled equations, we can transform the system of differential equations into a linear algebraic system (system of linear equations), due to the linearity of the equations and their constant coefficients. At this stage, the possible wave vector $k^{\mu}$ can be determined by the condition that the determinant of the coefficient matrix is zero, yielding the dispersion relation of gravitational waves. For the corresponding $k^{\mu}$, solving the linear system provides the relationship between gauge invariant perturbations and determines the polarization modes of gravitational waves. Additionally, due to spatial rotational symmetry, we can assume that gravitational waves propagate in the +z+z direction without loss of generality, simplifying the equations to be solved. Henceforth, we assume gravitational waves propagate along the +z+z direction.
@@
According to Eq. (IV), only $h^{TT}_{ij}$ contributes to the polarization modes. However, in the solutions corresponding to $m_{2}$ and $m_{3}$ in Eq. (117), one can obtian $h^{TT}_{ij}=0$, which means these solutions do not excite tensor mode gravitational waves. Therefore, tensor modes, specifically the ++ and $\times$ modes, can only propagate as massless modes at the speed of light or propagate with mass $m_{4}$. It is important to note that tensor mode gravitational waves, propagating at the speed of light, always exist.
@@ D.4 Case 4: no constraint variables
Here, the expressions for $\mathcal{A}$, $\mathcal{B}$, $\mathcal{C}$, and $\mathcal{D}$ are exceptionally complex, so they will not be listed here. Similarly, in the solutions corresponding to $m_{2}$ and $m_{3}$, $h^{TT}_{ij}=0$, so it does not actually contribute to tensor mode gravitational waves. The ++ and $\times$ modes that propagate at the speed of light always exist. The difference between the vector and scalar modes and Case 3 is that they possess more degrees of freedom. For the same reasons, a general analysis remains challenging.
@@ Appendix E Decoupled equations
$-\left(2A_{(1)}+2B_{(1)}+\frac{1}{2}C_{(1)}+E_{(1)}\right)N_{i}-\left(2B_{(1)}+4E_{(1)}\right)h_{i}-\left(8B_{(1)}+6E_{(1)}\right)q_{i}$ (130)
@@
$\left(A_{(1)}+\frac{1}{2}C_{(1)}\right)C_{i}-\left(A_{(1)}+\frac{1}{2}C_{(1)}\right)\bar{C}_{i}-\frac{1}{2}C_{(1)}K_{i}-\frac{1}{2}C_{(1)}\left(K_{i}-\Xi_{i}\right)$
@@
$\left.+2B_{(2)}\partial_{0}h_{i}+2B_{(2)}\partial_{0}N_{i}+8B_{(2)}\partial_{0}q_{i}+B_{(2)}\partial_{0}\Delta f_{i}\right)=0,$ (134)